OpenStudy (lacris):

I dont understand this, please help me understand :( I will give you a shiny medal :)

OpenStudy (lacris):

\[P(E and F) = P(E)0.45 and P(F) = 0.5\]

OpenStudy (lacris):

so I know its the conditional probability rule?

OpenStudy (lacris):

so the question is, what is P(F l E)?

OpenStudy (lacris):

do I multiply them?

MsBrains (ms-brains):

I want a shiny medal, but I don't understand this either. *^*

OpenStudy (lacris):

Oh awwie, xD

OpenStudy (lacris):

Do you know anyone who understands probability?

OpenStudy (rcp031):

RULE of Probability Probability OF failure+ Probability of success=1

OpenStudy (lacris):

so.. its 0.48+0.5 = 0.98?

OpenStudy (lacris):

The answer is supposed to be 0.96.. :(

OpenStudy (lacris):

What does the term (E | F) mean? Was I supposed to do: 0.98 *0.48* 0.5? If so why?

OpenStudy (agent0smith):

Numbers at the top aren't legible. P(E|F) means probability of E, given F already occurred. http://www.stat.yale.edu/Courses/1997-98/101/cond.gif

OpenStudy (agent0smith):

\[\large P(E|F) =\frac{ P( E \text { and } F)}{ P(F) }\]

OpenStudy (lacris):

Sorry, I had to do something and go somewhere, but Ahh! I see, thank you very much for explaining it to me.

OpenStudy (lacris):

My apologies for replying so late, but I see now what you mean. so since it's the "probability of E given F already occurred" Then that means, 0.48 / 0.5 = 0.96 since its just P(E and F) = 0.48 and P(E) = 0.5 so just: 0.48/0.5 as the equation you inserted! :D Thank you so much for taking your time to explain :D